# Immunity to Increasing Condition Numbers of Linear Superiorization versus Linear Programming

### v1

@@ Abstract
Given a family of linear constraints and a linear objective function one can consider whether to apply a Linear Programming (LP) algorithm or use a Linear Superiorization (LinSup) algorithm on this data. In the LP methodology one aims at finding a point that fulfills the constraints and has the minimal value of the objective function over these constraints. The Linear Superiorization approach considers the same data as linear programming problems but instead of attempting to solve those with linear optimization methods it employs perturbation resilient feasibility-seeking algorithms and steers them toward feasible points with reduced (not necessarily minimal) objective function values. Previous studies compared LP and LinSup in terms of their respective outputs and the resources they use. In this paper we investigate these two approaches in terms of their sensitivity to condition numbers of the system of linear constraints. Condition numbers are a measure for the impact of deviations in the input data on the output of a problem and, in particular, they describe the factor of error propagation when given wrong or erroneous data. Therefore, the ability of LP and LinSup to cope with increased condition numbers, thus with ill-posed problems, is an important matter to consider which was not studied until now. We investigate experimentally the advantages and disadvantages of both LP and LinSup on examplary problems of linear programming with multiple condition numbers and different problem dimensions.
Keywords: Condition number, ill-posed problem, Linear Programming, Linear Superiorization, bounded perturbations, feasibility seeking, Scipy, Gurobi.
@@ 1 Introduction
In this paper we investigate and discuss the behavior of the linear superiorization (LinSup) methodology for various different condition numbers and compare it to the behavior of multiple linear optimization solvers. We set up differently conditioned linear optimization problems in a manner that keeps them structurally similar in order to keep them comparable to each other. Then we solve these problems with the introduced superiorization methodology and with established implementations of the simplex algorithm and an interior point method. We compare the necessary runtimes in problems with multiple dimensions and discuss the observed effects. We show experimentally that superiorization is rather robust in comparison to scipy’s simplex implementations observing both faster runtimes and better objective function values.
@@ 1.1 Context and previous work
The superiorization methodology (SM) is a relatively new method that aims to improve an existing iterative algorithm by interlacing into it perturbation steps ([10]). In the context of optimization problems, this concept can be applied by interlacing into a feasibility-seeking algorithm (for example a projection method) perturbations of negative gradient steps that reduce the objective function values. Since its development, the SM has successfully been applied in various practical applications such as intensity-modulated radio therapy ([3]), image reconstruction ([11]) and telecommunication networks ([17]). In the particular context of linear optimization problems (LPs), there is a vast literature for both the feasibility problem (for example the Agmon-Motzkin-Schönberg algorithm [1, 24],[14, algorithm 5.4.1]) and the optimization problem (e. g., simplex algorithm, ellipsoid method [25]). Furthermore, for the LinSup case, the “guarantee problem of the superiorization method” has been answered positively. This problem is the question whether superiorization can actually converge to a feasbile point with a lower or equal objective function value than the point of the unsuperiorized version of the same feasibility-seeking algorithm. In [13] the authors employ the principle of concentration of measure to show the result with high probability. A similar conclusion for the nonlinear case is yet to be found.
This paper aims to compare the superiorization methodology with optimization algorithms for LPs in an organized, reproducible and fair manner for problems of varying difficulty, indicated by the problems’ dimensions and their condition numbers. Accordingly, our work is an extension of the results in [9], where a superiorization method for linear problems (called LinSup) was compared to the simplex method. We augment these results by varying the condition numbers across multiple problem instances. The condition number has been known to have a large impact on the performance of certain methods and often leads to the failure of an algorithm ([28]). As many real-world problems have an inherently high condition number, immunity against ill-conditioning of problems is a desirable property for any algorithm. Often it is enough to have a rough understanding of the order of magnitude of the condition number and there are efficient algorithms for its estimation ([15],[22]). There is a vast literature on preconditioning of matrix problems, which has been reviewed in [5].
@@ 1.2 Outline
We give a brief overview of the superiorization method and of condition numbers in Sections 2.1 and 2.2, respectively. We set up multiple linear optimization problems with varying condition numbers in Section 3.1 and run the superiorization method and several LP algorithms on them.
For details of our implementation of the superiorization method see Section 3.2. In Section 4 we present the results, followed by a discussion of remaining challenges and future work in Section 5.
2 Preliminaries
@@ 2.1 The superiorization methodology
The Superiorization Methodology (SM) developed from the investigation of feasibility-seeking models of some important real-world problems such as image reconstruction from projections and radiation therapy treatment planning. Feasibility-seeking algorithms, mainly projection methods, generate iterative sequences that (under reasonable conditions) converge to a point in the feasible set. Their main advantage is that they perform projections onto the individual sets whose intersection is the feasible set and not directly onto the feasible set and the underlying situation is that such projections onto the individual sets are more manageable.
When one wishes to find feasible points with a reduced, not necessarily minimized, value of an imposed objective function then the SM comes into play. The principle of the SM is to interlace into the iterates of a feasibility-seeking iterative process perturbations that will steer the iterates toward superior (meaning smaller or equal) objective function values without losing the overall convergence of the perturbed iterates to a feasible point. To this end “bounded perturbations” are used.
How all this is done is rigorously described in earlier papers on the SM, consult, e.g., [10] for a recent review, read also [20]. A key feature of the SM is that it does not aim for a constrained optimal function value, but is content with settling for a feasible point with reduced objective function value – reduced in comparison to the objective function value of a feasible point that would be reached by the same feasibility-seeking algorithm without perturbations. For many applications this is sufficient, in particular, whenever the introduction of an objective function is only a secondary goal. Fulfillment of the constraints, in this context, is considered by the modeler of the real-world problem to be much more important, see, e.g., [10, 11, 3].
Many papers on the SM are cited in [8] which is a Webpage dedicated to superiorization and perturbation resilience of algorithms that contains a continuously updated bibliography on the subject. This Webpage is a source for the wealth of work done in this field to date, including two special issues of journals [12] and [18] dedicated to research of the SM. Recent work includes [7, 19, 21, 23, 29].
@@ 2.2 Condition numbers
The relative condition number is a measure for the impact of deviations in the input data on the output data of a problem. In particular, it describes the factor of error propagation when given wrong or erroneous data. Let the function $f\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ represent some mathematical problem and let $x\in\mathbb{R}^{n}$ be its input, where $\mathbb{R}^{n}$ stands for the Euclidean $n$-dimensional space. Denote with $\tilde{x}\in\mathbb{R}^{n}$ the disturbed input data. Then the relative condition number of the problem at the point $x$ is defined as (see [31, equation (12.4)])
@@
as long as $f(x)\neq 0$. Otherwise, it is $\kappa_{\text{rel}}=\infty$. In particular, the condition number is independent of a chosen numerical algorithm for solving the problem $f$, but the algorithms convergence speed may depend on the magnitude of the condition number (see [28, section 4]). In the following we are interested in the condition number of matrices. When $f(x)=Ax$ describes the problem of matrix multiplication, where $A\in\mathbb{R}^{m\times n}$ and with $\|\cdot\|=\|\cdot\|_{2}$, the above formula becomes
@@
As $f$ is differentiable and writing $\tilde{x}-x=hv$, for some unit vector $v$ and $h=\|\tilde{x}-x\|$ we get
@@
The term on the right is called the condition number of the matrix $A$ (see [31, equation (12.15)]), where $A^{-1}$ denotes the inverse or, if $A$ is non-square, the pseudo-inverse of $A$:
$\kappa(A)=\|A\|\|A^{-1}\|=\frac{\sigma_{\text{max}}}{\sigma_{\text{min}}},$ (4)
where $\sigma_{\text{max}}$ and $\sigma_{\text{min}}$ are the maximal and minimal singular values of $A$, respectively. The condition number plays a significant role in the analysis of numerical problems and is subject to extensive studies in the literature ([16], [31], [15], [28]). Several methods exist to improve high condition numbers of ill-conditioned problems (these are the, so-called, pre-conditioning methods, see e. g., [5]) in order to increase the accuracy of calculated solutions. This is often necessary because many real-world applications give rise to condition numbers of significant magnitude. This is the key motivation for the investigation in this paper.
@@ 3.1 The problem formulation
We consider linear problems of the form
@@
$\kappa(A)=\frac{\sigma_{\text{max}}}{\sigma_{\text{min}}}=\frac{\sigma_{1}}{\sigma_{q}}.$ (6)
We want to construct a sequence of matrices $A_{\kappa}$ of a specified condition number $\kappa$ in such a way that they remain structurally similar to each other. To this end we reverse the singular value decomposition, that is, we create exactly one pair of $U$ and $V$ which contains the structure of the problem and construct, for different values of $\kappa$, diagonal matrices $\Sigma_{\kappa}$, which impose the condition number of the problem. Then, we calculate $A_{\kappa}:=U\Sigma_{\kappa}V$. As any matrix has a singular value decomposition, this makes it possible to define any matrix via this approach too. In our construction we focus on matrices of full rank, because otherwise one can remove rows or columns until full rank is achieved. In order to construct $\Sigma$, setup $q=\min(m,n)=\operatorname{rank}(A)$ singular values as
$\sigma_{i}=\frac{t}{z_{i}}+\frac{1-t}{s}$ (7)
where $t=\frac{\kappa-1}{q-1}$, $s=10$ and $z_{i}=\frac{si}{q}$. Note that this choice is somewhat arbitrary. We chose this setup of the $\sigma_{i}$, because in real-world applications the singular values often behave approximately proportional to $\frac{1}{i}$ (instead of linear which seems like an obvious first choice for our problem). The parameter $s$ is used to control the magnitude of the singular values since in this model we always have $\sigma_{q}=\frac{1}{s}$. Due to the choice of $t$ it is easy to see that
@@
as desired. For $U$ and $V$ choose random semi-orthogonal matrices and set $A=U\cdot\operatorname{diag}(\sigma_{1},\dots,\sigma_{q})\cdot V$. Furthermore, set $b=A\mathbbm{1}+\mathbbm{1}$, $u=-\ell=100\cdot\mathbbm{1}$ and randomly choose $c$. This choice of parameters implies the feasibility of $x=\mathbbm{1}$. We set up this problem for multiple dimensions, ranging from $80\times 100$ to $2000\times 2500$ (cf. [9]).
@@ 3.2 The superiorization algorithm
We aim to apply separately linear programming and the superiorization methodology for the data $A=(a^{i})_{i=1}^{m},b,c,\ell,u$ that appears in (5). In the SM we chose for the feasibility-seeking algorithm the projection method of Agmon Motzkin and Schönberg (AMS) with overshoot $0<r:=10^{-3}$ as the “basic algorithm”. This algorithm cyclically projects onto the individual half-spaces $\langle a^{i},x\rangle\leq b_{i}$ via (as defined in [27, p. 411])
$\mathcal{A}_{i}(x):=\begin{cases}x-\frac{\langle a^{i},x\rangle-b_{i}+r\|a^{i}\|}{\|a^{i}\|^{2}}a^{i}\text{, if $\langle a^{i},x\rangle>b_{i}$}\\ x\text{, otherwise}\end{cases}$ (9)
@@
$\mathcal{A}:=\mathcal{A}_{m}\circ\dots\circ\mathcal{A}_{1}.$ (10)
The parameter $r$ describes, how far into the half-space the projection maps (i. e., beyond the bounding hyperplane). For the direction vectors in the perturbations used in the SM we chose the normalized negative gradient of the objective function in (5), which is constant throughout and equals $-\frac{c}{\|c\|}$. For the step-sizes $\eta_{k}$ we take an exponentially decreasing null sequence $(10\cdot 0.99^{n})_{n\in\mathbb{N}}$ with resets every $\tau_{\text{reset}}=20$ iterations as described in [2, p. 6]. The starting step-size $\eta_{0}=10$ is decreased by the kernel $\alpha=0.99$ in each iteration, unless there is a reset. In that case, set $\eta_{k}=\eta_{0}\alpha^{\rho}$, where $\rho$ is the number of previous resets during this run.
In other words, the $k$-th iteration consists of a gradient step $-\eta_{k}\frac{c}{\|c\|}$ followed by a cyclic sweep of projections onto the half-spaces via $\mathcal{A}$ as given in (9) and (10). This process was repeated, until the iterate $x^{k}$ was feasible up to a tolerance of $\epsilon=10^{-8}$ and the relative change $\frac{\|x^{k}-x^{k-1}\|}{\|x^{k-1}\|}$ from the previous iterate was negligible, i. e., became smaller than $\epsilon$. Algorithm 1 contains the pseudocode for the described process. All runs were initialized at the all-ones vector $x^{0}=\mathbbm{1}$. The parameters $\epsilon$, $\alpha$, $\eta_{0}$, $r$, $\tau_{\text{reset}}$ can be adjusted for individual preferences or a particular problem.
Data: $A$, $b$, $c$, $\ell$, $u$, $x^{0}$
@@
Parameters : $\epsilon=10^{-8}$, $\alpha=0.99$, $\eta_{0}=10$, $r=10^{-3}$, $\tau_{\text{reset}}=20$

### v2

@@ Abstract
Given a family of linear constraints and a linear objective function one can consider whether to apply a Linear Programming (LP) algorithm or use a Linear Superiorization (LinSup) algorithm on this data. In the LP methodology one aims at finding an optimal point, i.e., a point that fulfills the constraints and has the minimal value of the objective function over these constraints.
The Linear Superiorization approach considers the same data as linear programming problems but instead of attempting to solve those with linear programming methods it employs perturbation resilient feasibility-seeking algorithms and steers them toward a feasible point with reduced (not necessarily minimal) objective function value. This aim of the superiorization method (SM) is less demanding than aiming to reach full-fledged constrained optimality and it places more importance on reaching feasibility than on reaching optimality.
Previous studies (e. g. [12]) compared LP and LinSup in terms of their respective outputs and the resources they use. This paper is a follow-up analysis of [12], where we investigate classical LP approaches and LinSup in terms of their sensitivity to condition numbers of the system of linear constraints.
Condition numbers are a measure for the impact of deviations in the input data on the output of a problem and, in particular, they describe the factor of error propagation when given wrong or erroneous data.
Therefore, the ability of LP and LinSup to cope with increased condition numbers, thus with ill-posed problems, is an important matter to consider which was not studied until now. We investigate experimentally the advantages and disadvantages of both LP and LinSup on exemplary problems of linear programming with multiple condition numbers and different problem dimensions.
Keywords: Condition number, ill-posed problem, Linear Programming, Linear Superiorization, bounded perturbations, feasibility-seeking, Scipy, Gurobi.
@@ 1 Introduction
In this paper we investigate experimentally and discuss the behavior of the linear superiorization (LinSup) methodology for various different condition numbers and compare it to the behavior of several linear programming (LP) solvers. The paper is in the spirit of [12] and contains a follow-up analysis regarding the impact of condition numbers on the runtime of the respective algorithms.
We set up linear programming problems with different condition numbers of the linear constraints in a manner that keeps the problems structurally similar in order to enable the comparative experimental study. Then we apply to these problems the LinSup superiorization method and some established implementations of the simplex algorithm and an interior point method.
Imposing stopping rules that are fair toward all algorithms, we compare the outputs on problems with increasing condition numbers and on problems with increasing dimensions.
1.1 The experimental methodology
LinSup is not an LP solver. As an offspring of the superiorization methodology (SM), LinSup perturbs linear feasibility-seeking algorithms to reduce objective function values while retaining the feasibility-seeking nature of the algorithm and without paying a high computational price.
The declared aim of LinSup is to find a feasible point whose objective function value is “superior” (meaning smaller or equal) to that of a feasible point that can be reached by the corresponding unperturbed feasibility-seeking algorithm, i.e., the exactly same feasibility-seeking algorithm that the LinSup employs. This aim is less demanding than full-fledged linear programming but more demanding than plain feasibility-seeking. [^1] Although we came across it after the SM was developed, one can borrow support for the reasoning of the SM from the American scientist and Nobel-laureate Herbert Simon who was in favor of “satisficing” rather than “maximizing”. Satisficing is a decision-making strategy that aims for a satisfactory or adequate result, rather than the optimal solution. This is because aiming for the optimal solution may necessitate needless expenditure of time, energy and resources. The term “satisfice” was coined by Herbert Simon in 1956 ([33]). See also: https://en.wikipedia.org/wiki/Satisficing.
In view of this difference between LinSup and an LP solver, but recognizing that both are initialized at a starting point which is not known to be a feasible point, we maintain the following experimental methodology. For a given problem we run LinSup until it reaches some predetermined fixed small threshold level of “infeasibility” (i.e., “constraints violation”). When LinSup stopped according to this stopping rule we record the objective function value at this point and the run-time elapsed.
Then we run on the same problem, starting from the same initialization point, the LP solver that we wish to include in our comparative study, and stop it when it reaches the same predetermined fixed level of “infeasibility” that was used for stopping the LinSup run. When stopped we record the objective function value at this point and the run-time elapsed for this LP solver as well.
We also record the infeasibility and the objective function values along the whole run-time and use these in order to demonstrate the behavior of each algorithm by plotting its objective function values and its infeasibility values versus time, respectively. More details are supplied later in the paper.
@@ 1.2 Context and previous work
The superiorization methodology (SM) is a relatively new method that aims to improve the performance of an existing iterative algorithm by interlacing into it perturbation steps ([13]). In the context of optimization problems, this concept can be applied by interlacing into a feasibility-seeking algorithm (for example a projection method) perturbations of negative gradient steps that reduce the objective function values.
Since its development, the SM has successfully been applied in various practical applications such as intensity-modulated radio therapy ([4]), image reconstruction ([14]) and telecommunication networks ([20]). In the particular context of linear programming problems (LPs), there is a vast literature for both feasibility-seeking problems (for example the Agmon-Motzkin-Schönberg (AMS) algorithm [2, 28],[17, algorithm 5.4.1]) and for the LP problem (e.g., simplex algorithm, ellipsoid method [29]).
Furthermore, for the LinSup case, the “guarantee question of the SM” has been answered positively. This is the question whether superiorization can actually converge to a feasible point with a smaller or equal objective function value than that of the point reached by the un-superiorized version of the same feasibility-seeking algorithm. In [16] the authors employ the principle of concentration of measure to show that this result holds with high probability. A similar conclusion for the nonlinear case is yet to be found.
In this paper we compare experimentally the SM with optimization algorithms for LPs in an organized, reproducible and fair manner for problems of varying difficulty, indicated by the problems’ dimensions and their condition numbers.
Accordingly, our work is an extension of the results in [12], where LinSup was compared to the simplex method. We augment these results by varying the condition numbers across multiple problem instances. The condition number of the linear system has been known to have a significant impact on the performance of certain methods, often leading to the failure of an algorithm ([31]). Since many real-world problems have an inherently high condition number, immunity against ill-conditioning of problems is a desirable property for any algorithm. Often it is enough to have a rough understanding of the order of magnitude of the condition number and there are efficient algorithms for its estimation ([18],[25]). For literature on preconditioning of matrix problems see, e.g., [8].
1.3 Our results
In this project we used LP data aiming at the task of finding a feasible point with reduced (not necessarily minimal) objective function value. In our numerical work the superiorization method LinSup performed better than LP solvers for this task. This better performance, for this task, was investigated for problems with increasing condition numbers of the linear systems and with increasing problem sizes. We found that in each of these cases LinSup delivered better results than some established LP solvers.
@@ 1.4 Outline
We give a brief overview of the superiorization method and of condition numbers in Sections 2.1 and 2.2, respectively. We set up multiple linear programming problems with varying condition numbers in Section 3.1 and run the LinSup method and several LP algorithms on them.
For details of our implementation of the LinSup method see Section 3.2. We give a brief description of the optimization suites that we use for our comparisons in Section 3.3 and of the experimental setup in Section 3.4. In Section 4 we present the numerical results, followed by a discussion of remaining challenges and future work in Section 5.
2 Preliminary background
@@ 2.1 The superiorization methodology
The Superiorization Methodology (SM) evolved from the investigation of feasibility-seeking models of some important real-world problems such as image reconstruction from projections and radiation therapy treatment planning.
Feasibility-seeking algorithms, mainly projection methods, generate iterative sequences that (under reasonable conditions) converge to a point in the feasible set. Their main advantage is that they perform projections onto the individual sets whose intersection is the feasible set and not directly onto the feasible set and the underlying premise is that such projections onto the individual sets are more manageable.
When one wishes to find feasible points with a reduced, not necessarily minimized, value of an imposed objective function then the SM comes into play. The principle of the SM is to interlace into the iterates of a feasibility-seeking iterative process perturbations that steer the iterates toward superior (meaning smaller or equal) objective function values without losing the overall convergence of the sequence of perturbed iterates to a feasible point. To this end “bounded perturbations” are used.
How all this is done is rigorously described in earlier papers on the SM, consult, e.g., [13] for a recent review, read also [23]. A key feature of the SM is that it does not aim for a constrained optimal function value, but is content with settling for a feasible point with reduced objective function value – reduced in comparison to the objective function value of a feasible point that would be reached by the same feasibility-seeking algorithm without perturbations.
For many applications this is sufficient, in particular, whenever the introduction of an objective function is only a secondary goal. Fulfillment of the constraints, in this context, is considered by the modeler of the real-world problem to be much more important, see, e.g., [13, 14, 4].
Many papers on the SM are cited in [11] which is a Webpage dedicated to superiorization and perturbation resilience of algorithms that contains a continuously updated bibliography on the subject. This Webpage (which includes 199 entries as of December 10, 2025) is a source for the wealth of work done in this field to date, including two special issues of journals [15] and [21] dedicated to research of the SM. Recent work includes [10, 1, 22, 24, 27, 32].
We find especially interesting the recent work of Ma et al. [26] who proposed a novel decomposition framework for derivative-free optimization (DFO) algorithms which significantly extended the scope of current DFO solvers to larger-scale problems. They proved that their proposed framework closely relates to the superiorization methodology.
Information about feasibility-seeking algorithms and the convex feasibility problem, that lie at the foundation of the SM approach, can be found, e.g., in Bauschke and Borwein’s SIAM review paper [7] and in [6].
@@ 2.2 Condition numbers
The relative condition number is a measure of the impact of deviations in the input data on the output data of a problem. In particular, it describes the factor of error propagation when given wrong or erroneous data.
Let the function $f\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ represent some mathematical problem and let $x\in\mathbb{R}^{n}$ be its input, where $\mathbb{R}^{n}$ stands for the Euclidean $n$-dimensional space. Denote with $\tilde{x}\in\mathbb{R}^{n}$ the disturbed input data. Then the relative condition number of the problem at the point $x$ is defined as (see [34, equation (12.4)])
@@
as long as $f(x)\neq 0$. Otherwise, it is $\kappa_{\text{rel}}=\infty$.
In particular, the condition number is independent of a chosen numerical algorithm for solving the problem $f$, but the algorithms convergence speed may depend on the magnitude of the condition number (see [31, section 4]).
In the following we are interested in the condition number of matrices. When $f(x)=Ax$ describes the problem of matrix multiplication, where $A\in\mathbb{R}^{m\times n}$ and with $\|\cdot\|=\|\cdot\|_{2}$, the above formula becomes
@@
Since $f$ is differentiable and by writing $\tilde{x}-x=hv$, for some unit vector $v$ and $h=\|\tilde{x}-x\|$ we get
@@
The term on the right is called the “condition number of the matrix AA” (see [34, Equation (12.15)]), where $A^{-1}$ denotes the inverse or, if $A$ is non-square, the pseudo-inverse of $A$,
$\kappa(A):=\|A\|\|A^{-1}\|=\frac{\sigma_{\text{max}}}{\sigma_{\text{min}}},$ (4)
where $\sigma_{\text{max}}$ and $\sigma_{\text{min}}$ are the maximal and minimal singular values of $A$, respectively.
The condition number plays a significant role in the analysis of numerical problems and is subject to extensive studies in the literature ([19], [34], [18], [31]). Several methods exist to improve high condition numbers of ill-conditioned problems (these are the, so-called, pre-conditioning methods, see e.g., [8]) in order to increase the accuracy of calculated solutions. This is often necessary because many real-world applications give rise to condition numbers of significant magnitude. This is the key motivation for the investigation in this paper.
@@ 3.1 The problem formulation
We consider LP problems of the form
@@
$\kappa(A):=\frac{\sigma_{\text{max}}}{\sigma_{\text{min}}}=\frac{\sigma_{1}}{\sigma_{q}}.$ (6)
We want to construct a sequence of matrices $A_{\kappa}$ of a specified condition number $\kappa$ in such a way that they remain structurally similar to each other. To this end we reverse the singular value decomposition, that is, we create exactly one pair of $U$ and $V$ which contains the structure of the problem and construct, for different values of $\kappa$, diagonal matrices $\Sigma_{\kappa}$, which impose the condition number of the problem. Then, we calculate $A_{\kappa}:=U\Sigma_{\kappa}V$.
As any matrix has a singular value decomposition, this makes it possible to define any matrix via this approach too. In our construction we focus, without loss of generality, on matrices of full rank, because otherwise one can remove rows or columns until full rank is achieved. For any chosen value of $\kappa$ we construct the diagonal matrix $\Sigma_{\kappa}$, by setting $q:=\min(m,n)=\operatorname{rank}(A)$ and defining singular values as
$\sigma_{i}:=\frac{t}{z_{i}}+\frac{1-t}{s},$ (7)
where $t:=\frac{\kappa-1}{q-1}$, $s=10$ and $z_{i}=\frac{si}{q}$. We chose this setup of the singular values $\sigma_{i}$, because in real-world applications the singular values often behave approximately proportional to $\frac{1}{i}$ (instead of linear which seems like an obvious first choice for our problem).
The parameter $s$ is used to control the magnitude of the singular values since in this model we always have $\sigma_{q}=\frac{1}{s}$. Due to the choice of $t$ it is easy to see that
@@
as desired. For $U$ and $V$ we generate random semi-orthogonal matrices and set $A=U\cdot\operatorname{diag}(\sigma_{1},\dots,\sigma_{q})\cdot V$.
Furthermore, in order to guarantee the non-emptiness of the feasible set in (5) we set $b:=A\mathbbm{1}+\mathbbm{1}$, $u=-\ell=100\cdot\mathbbm{1}$ and randomly choose $c$. This choice of parameters implies the feasibility of $x=\mathbbm{1}$.
We use the above set up to generate LP problems for multiple dimensions, with matrix sizes ranging from $80\times 100$ to $4000\times 5000$ (cf. [12]).
@@ 3.2 The superiorization algorithm
We aim to apply separately LP solvers and the LinSup method for the data $A=(a^{i})_{i=1}^{m},b,c,\ell,u$ that appears in (5). In LinSup, presented here in Algorithm 1, we chose for the feasibility-seeking algorithm the projection method of Agmon Motzkin and Schönberg (AMS) as the “basic algorithm”.
This algorithm sequentially cyclically projects onto the individual half-spaces $\langle a^{i},x\rangle\leq b_{i}$ using (see, e.g., [12, p. 6]) for each individual projection the closed-form formula
$P_{i}(x):=\begin{cases}x-\lambda\frac{\langle a^{i},x\rangle-b_{i}+r\|a^{i}\|}{\|a^{i}\|^{2}}a^{i}\text{, if $\langle a^{i},x\rangle>b_{i},$}\\ x\text{, otherwise,}\end{cases}$ (9)
@@
$\mathcal{A}:=P_{m}\circ\dots\circ P_{1}.$ (10)
The parameter $\lambda$ describes a relative overshoot or undershoot of the projection. For our purposes we set $\lambda=1$ so that the projection $P_{i}$ maps precisely onto the $i$-th individual half-space $\langle a^{i},x\rangle\leq b_{i}$.
For the direction vectors in the perturbations used in LinSup of Algorithm 1 we chose the normalized negative gradient of the objective function in (5), which is constant throughout and equals $\frac{c}{\|c\|}$.
For the step-sizes $\eta_{k}$ we take an exponentially decreasing null sequence $(10\cdot 0.99^{n})_{n\in\mathbb{N}}$ with restarts every $\tau_{\text{restart}}=20$ iterations as described in [3, p. 6]. The starting step-size $\eta_{0}=10$ is decreased by the kernel $\alpha=0.99$ in each iteration, unless there is a restart. In that case, we set $\eta_{k}=\eta_{0}\alpha^{\rho}$, where $\rho$ is the number of previous restarts during this run.
In other words, the $k$-th iteration consists of a gradient step $\eta_{k}\frac{c}{\|c\|}$ followed by a cyclic sweep of projections onto the half-spaces via $\mathcal{A}$ as given in (9) and (10).
This process was repeated, until the iterate $x^{k}$ became feasible up to the infeasibility tolerance of $\varepsilon=10^{-8}$ and the relative change $\frac{\|x^{k}-x^{k-1}\|}{\|x^{k-1}\|}$ from the previous iterate was negligible, i.e., became smaller than $\varepsilon$. Algorithm 1 is the pseudocode for this LinSup process. All runs were initialized at the all-zeros vector $x^{0}=0$. The parameters $\varepsilon$, $\alpha$, $\eta_{0}$, $\lambda$, $\tau_{\text{restart}}$ can be adjusted for individual preferences or a particular problem.
Data: AA, $b$, $c$, $\ell$, $u$, $x^{0}$
@@
Parameters : $\varepsilon=10^{-8}$, $\alpha=0.99$, $\eta_{0}=10$, $\lambda=1$, $\tau_{\text{restart}}=20$